# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ VIII.2 Time-crystal behavior
On top of being globally ordered and time-dependent, the 3D swap phase possesses another notable property: its coherence time diverges because of its many-body nature, making it a time crystal [128, 134, 135, 136, 129, 137, 138].
To understand this feature, consider a single limit cycle oscillator subject to noise. This can be represented by the zero-dimensional MF equation of the NR Ising model with added noise, namely
$\partial_{t}m_{\alpha}=-m_{\alpha}+\tanh(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta})+\eta_{\alpha}\;,$ (54)
@@
$C_{x}\left(\tau\right)=\frac{\langle x\left(t\right)x\left(t+\tau\right)\rangle_{t,\Omega}}{\langle x\left(t\right)^{2}\rangle_{t,\Omega}}.$ (55)
Note that $\langle x\rangle=0$. If this was not the case, one should replace $x(t)$ with $x(t)-\langle x\rangle$ in Eq. (55). As seen in Figure 18, even though the magnetization performs quit regular oscillations, its auto-correlation function decays, thereby indicating that this oscillator eventually forgets its initial phase [171, 172, 124, 173, 174, 175, 176, 177].
In many-body systems, however, interactions can overcome noise, producing coherent oscillations over arbitrarily long periods [122, 95, 98, 108, 178, 110, 112, 94, 93, 111, 179, 106, 107, 110, 109, 180, 178, 96, 181, 182, 91, 90]. Figure 18 shows that in the 3D NR Ising model, temporal correlations persist and do not decay over large times (panels b-c and e-f), making it a robust clock. We stress that this robust clock is an emergent feature of our model which does not rely on any periodic drive externally imposing a phase.
The different behaviors of the time correlations in the noisy clock Eq. (54) and in the 3D NR Ising can be respectively compared to spatial density correlations in a liquid (which decay with distance) and a solid (which is maintained over arbitrarily large distances). Such comparison makes (i) a “time liquid” and (ii) a “time crystal” which has infinite coherence time.
We find the coherence time of the swap phase in the 3D NR Ising model by determining the decay rate of the auto-correlation’s envelope, $\tilde{C}_{M_{A}}(\tau)$. We construct this envelope $\tilde{C}_{M_{A}}(\tau)$ by interpolating the peaks of $|C_{M_{A}}(\tau)|$. We then implicitly define the coherence time $\tau_{c}$ as the time for which
@@
i.e. linearly with the number of spins, thereby proving the perfect time-crystal behavior of the 3D NR Ising in the thermodynamic limit. Note that a similar scaling for $\tau_{c}$ has been observed in a periodically driven 2D system in Ref. [183].
We can combine Eq. (57) with the conjecture of Ref. [125] to obtain a bound on the minimum rate of entropy production required to maintain the oscillations. This conjecture bounds the average entropy per oscillation $\Delta S$ by the number of coherent oscillations $\mathcal{N}\equiv\tau_{c}/T_{\rm osc}$ through the inequality $\Delta S\geq 4\pi^{2}\mathcal{N}$. As a consequence, the rate of entropy production per unit volume $\dot{s}\equiv\Delta S/(L^{d}T_{\rm osc})$ must satisfy
@@
Note that in the stable swap phase $T_{\rm osc}$ does not depend on $L$ (it converges to a constant as $L\to\infty$, see Fig. 26a in Sec. IX.1). As $\tau_{c}(L)\sim L^{d}$, this bound converges to a finite value when $L\to\infty$, meaning that there is a finite cost per oscillator to maintain the time-crystal state. Note that the fit in inset of Fig. 19 provides the prefactor for the example considered there.
Fig. 18: Time correlations in the NR Ising model. (a) Time evolution of $m_{A}$ in the zero-dimensional MF equation with added noise. (b) Time evolution of a single spin $\sigma^{A}_{i}$ in the 3D NR Ising model. (c) Time evolution of the total magnetization of $A$-spins, $M_{A}$, in the 3D NR Ising model. (d-f) Auto-correlation function of $m_{A}$, $\sigma^{A}_{i}$, and $M_{A}$, respectively (defined in Eq. (55)). Panels a and d are obtained from numerical integration of Eq. (54) with the parameters: $\tilde{J}=1.2$, $\tilde{K}=0.12$, $\zeta=0.0009$, $dt=0.01$. Panels b-c and e-f are obtained from MC simulation with the parameters: $\tilde{J}=1.6$, $\tilde{K}=0.1$, $L=80$.
Fig. 19: Diverging coherence time of the swap phase. The envelope $\tilde{C}_{M_{A}}(\tau)$ of the auto-correlation function of the total magnetization of $A$-spins in the 3D NR Ising model for linear system sizes $L=12,16,20,24,28,32$ (from bottom to top). Coupling parameters $\tilde{J}=1.6$ and $\tilde{K}=0.1$. The coherence time $\tau_{c}$ (Eq. (56)) is calculated from the intersection of $\tilde{C}_{M_{A}}(\tau)$ with $e^{-1}$ (dashed black line). Inset: $\tau_{c}$ as a function of $L$ on a log-log scale. Dots are calculated from the data in the main panel and line is best linear fit. The coherence time is proportional to $L^{d}$, supporting a time-crystal behavior. Obtained from MC simulations.
Finally, we note that ordered oscillations and infinite coherence time are not decoupled properties. It is plausible that a locally coupled system can maintain ordered oscillations over arbitrarily large distances only if all its parts adhere to the same prescribed phase, determined by the initial conditions.
@@ VIII.3 The phase transition from disorder to swap in the high $(\tilde{J},\tilde{K})$ regime
Fig. 20: Destruction of the second-order phase transition due to scroll-waves in the high $\tilde{K}$ and high $\tilde{J}$ regime. (a) $R$ vs. $\tilde{K}$ for $\tilde{J}=2$. phase-space angular momentum is shown in the inset. (b) Snapshot of a 2D slice from a 3D simulation with the parameters $\tilde{J}=2$, $\tilde{K}=0.9$, $L=320$, marked by a black hollow circle on panel a. (c) The same snapshot coarse-grained over $10\times 10$ lattice sites. The snapshots show the destruction of order by scroll waves (appearing as spiral defects due to the 2D projection). Obtained from MC simulations.
@@
At high $\tilde{J}$ and $\tilde{K}$ values, the transition from disorder to swap is different than for small $\tilde{J}$ and $\tilde{K}$ values. This is seen in Fig. 12 where the change in $R$ and $\mathcal{L}$ at the phase transition appears to become more abrupt in the upper right region, compared to the lower left region. We further analyze this in Fig. 20a by plotting the order parameters as a function of $\tilde{K}$, for fixed $\tilde{J}=2$, and various system sizes. The plots of $R$ and $\mathcal{L}$ vs. $\tilde{K}$ do not seem to converge to a second-order phase transition in which the order parameters are continuous. Instead, an abrupt change in $R$ and $\mathcal{L}$ emerges at high system sizes, suggesting a discontinuity as in first-order phase transitions. A clear critical point cannot be identified since convergence is not reached at $L=320$, the largest linear system size in our simulations (see $0.8\lesssim\tilde{K}\lesssim 1.1$ region in Fig. 20a).
@@
The sharp decrease in synchronization is caused by the emergence of highly noisy scroll waves that destructs the order. This is seen in Fig. 20b showing a 2D snapshot taken from the 3D simulation. Scroll waves are the 3D analog of spiral waves, having line defects that are typically closed into rings instead of point defects [136, 184, 84, 185] (see also Movie 3 and Fig. 29 in Sec. X in which the scroll waves are more visible). They also have spiral cross sections, which is seen in the 2D-cut $\theta$ snapshot in Fig. 20b and in its coarse-grained version, Fig. 20c (the latter figure makes the defects more apparent by averaging out the noise). We discuss scroll waves in more detail in Sec. X.
@@
The destruction of the swap phase by scroll waves is slightly reminiscent of the destruction of order in 2D due to spiral defects (compare Fig. 10 with Fig. 20a), except that here, it appears that the destruction only occurs for a limited range of high $\tilde{J}$ and $\tilde{K}$ values whereas smaller values are protected. Due to the size limitation of our simulations we cannot prove what portion of the 3D swap phase is destructed by the scroll waves or what kind of phase transition is associated with this destruction, although it appears to be first-order-like (in the sense that the order parameters are discontinuous).
VIII.4 Why is the swap stable in 3D?
@@
Fig. 21: Continuous symmetry breaking of the swap phase. (a) Comparison of discrete and continuous symmetry breaking in equilibrium and out-of-equilibrium models. (b) $M_{B}$ vs. $M_{A}$ for $\tilde{K}=0.5$, $\tilde{J}=2$, and $L=80$. i,ii, and iii show increasing numbers of MC sweeps. Since the oscillation period is incommensurate, all phases are sampled in the infinite time limit. Hence, a continuous symmetry is broken. (c) Oscillation period of the magnetizations, $T_{\rm osc}$, in the 3D system as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and $L=80$. $T_{\rm osc}$ is measured as twice the average time between two subsequent sign flips of $M_{A}$. The continuous dependence of $T_{\rm osc}$ on the coupling parameter indicates incommensurate oscillations. (d-e) comparison between domain walls in the Ising model (first row) and in the NR Ising model (second row). At $t=0$, the two systems were at steady state and then perturbed by a domain wall. The domain wall is created by flipping the magnetization (Ising) and shifting the oscillation phase (NR Ising) for all spins located at $y>L/2$. System parameters are: $\tilde{J}=1.4$ (Ising) and $\tilde{J}=1.5$, $\tilde{K}=0.1$ (NR Ising), and $L=320$. i and iii are Snapshots of $\sigma$ (Ising) and $\theta$ (NR Ising, see color code in Fig. 2) in a 2D slice taken from a 3D simulation at $t=0$ and $t=1157$, respectively. ii Kymograph of $|{\rm d}{\bar{m}}/{\rm d}{y}|$ (Ising) and $|{\rm d}{\bar{\theta}}/{\rm d}{y}|$ (NR Ising), defined by the discrete symmetric derivative ${\rm d}f/{\rm d}y\equiv\left|\left(f_{y+5}-f_{y-5}\right)/10\right|$, where $\bar{m}$ is the magnetization and $\bar{\theta}$ is the phase, both averaged over the $xz$ plane. The smearing out of the dark pink region indicates the dissipation of the domain wall. Panels b-e are obtained from MC simulations.
@@
Our numerical evidence point to the existence of a stable swap phase. This relates to a longstanding question: can time-periodic states of noisy many-body systems with short-range interactions be stable in the thermodynamic limit [95, 98, 122, 108, 178, 110, 112, 94, 93, 111, 179, 106, 107, 110, 109, 180, 178, 96, 181, 182]. References [95, 98, 96] distinguished two cases, depending on whether the spontaneously broken time-translation invariance is discrete or continuous (Fig. 21a). In systems that evolve in discrete time as our MC simulation, this depends on the type of oscillations: commensurate vs. incommensurate. When the period of the oscillations is commensurate, i.e., a multiple of the discrete time-unite (a case referred to as “periodic $k$-cycle”), a discrete time-translation symmetry is broken. In that case, it was argued in [95] that the growth of out-of-phase droplets  [^4] Note that these potential droplets are not the droplets that destabilize the static-order phase (Fig. 23). Instead, these are droplets of a homogeneously oscillating phase, immersed in an oscillatory background destabilizes long-range ordered oscillations. That is because flat domain walls are expected to have a non-vanishing velocity (unlike equilibrium [187]), indicating that a fluctuation producing a large enough droplet, can expand rather than shrink (similar destabilization mechanism is observed in discrete-symmetry flocks [188]).
@@
In the incommensurate case, i.e. when the oscillation period is not a multiple of the discrete time-unite (e.g. the time step of the MC simulation), the discrete-time system is equivalent to continuous-time one (the discreteness is irrelevant). The oscillatory state then breaks a continuous time-translation symmetry, which can be seen by the fact that the oscillation’s phase at a given time can take any value between 00 and $2\pi$. In continuous symmetry breaking, domain walls smear out (as in the XY model) and thus the destabilization-by-droplets argument does not apply [98, 96].
@@
In the NR Ising model studied here, the oscillation period is incommensurate with the Monte-Carlo time step. This is seen both in Fig. 21b where all the points on the limit cycle are visited in the infinite-time limit, and in Fig. 21c where the period of the oscillation, measured as twice the average time between two subsequent sign flips of $M_{A}$, is shown to change continuously with the interaction parameters (a continuously changing variable is in general incommensurate with $1$). This means that, while stable oscillations are not guaranteed, they are not ruled out by the droplet argument in Ref. [95], see also Ref. [182]. Figure 21d-e further exemplifies this: it shows an analysis of a 3D simulation in which the system is perturbed from a stable swap state by introducing a flat domain wall. As time evolves, the domain wall is shown to widen and dissipate (on top of slightly moving in the $+\hat{y}$ direction), unlike a domain wall in the Ising model, supporting the picture of no steady domain-walls between oscillating states.
The broken continuous symmetry provides a reasoning for the resemblence of the NR Ising model with the XY ($O(2)$) model, observed both in the absence of time-dependent order in 2D and its stabilization in 3D with the corresponding critical exponents calculated in Sec. VIII.1. Moreover, it suggests an analogy with the Mermin-Wagner theorem [105, 81] in the time domain [104, 87]: continuous changes in the oscillation phase over space may behave as generalized Goldstone modes and their fluctuations preclude order in 2D yet allow for order in 3D and beyond.
@@
An analogy between (de)synchronization and surface roughening provides further reasons to expect the swap phase to be stable in 3D and not in 2D. If we add random noise to the MF equation (28), it takes the form of a noisy reaction-diffusion system
@@
where $\vec{X}=x_{1},..,x_{n}$ and $D$ can be a matrix. Now assume that the noiseless and diffusion-less version of Eq. (59) has a limit cycle in steady state. To test whether the addition of diffusion and noise allows for synchronized oscillations, we assume that we are deep within the synchronized regime, where fluctuations from the global limit cycle are small and can be treated perturbatively. In this case, we can focus on the phase variable that describes the position of the system on the limit cycle, while neglecting other degrees of freedom. This is known as a phase reduction, see Refs. [151, 189, 190, 191]. Then, Eq. (59) can be mapped to the known Kardar–Parisi–Zhang (KPZ) equation [151, 98, 101, 102],
$\frac{\partial\Psi}{\partial t}=\alpha\nabla^{2}\Psi+\beta\left(\nabla^{2}\Psi\right)+{\rm noise}$ (60)
where the deviation $\Psi$ of the local phase with respect to the phase of the deterministic limit cycle plays the role of the surface height in the KPZ equation (Fig. 22). The KPZ equation predicts that interfaces are always rough in 2D, but that there is a phase transition between rough and smooth interfaces in 3D [192]. This suggests, by analogy, that a coherent oscillating phase can exist in 3D, but not in 2D. Note, however, that the variable $\Psi$ in Eq. (60) is an angle (defined on a compact space, the circle $S^{1}$), so the results on the usual KPZ equation cannot be directly applied. In particular, the fact that the phase $\Psi$ is compact allows for the existence of topological defects, that may change the physics. We refer to Refs. [101, 102, 98, 193, 194, 195] for further details on the compact KPZ equation. In 3D, defect rings are expected to be small (at least for certain range of $\tilde{J}$ and $\tilde{K}$) and hence not to completely destroy long-range order [98].
@@
IX Destruction of static-order phase via droplets

### v2

@@ VIII.2 Time-crystal behavior
On top of being globally ordered and time-dependent, the 3D swap phase possesses another notable property: its coherence time diverges because of its many-body nature, making it a time crystal [134, 140, 141, 142, 135, 143, 144, 146, 147].
To understand this feature, consider a single limit cycle oscillator subject to noise. This can be represented by the zero-dimensional mean-field equation of the nonreciprocal Ising model with added noise, defined as
$\frac{{\rm d}m_{\alpha}}{{\rm d}t}=-m_{\alpha}+\tanh(\tilde{J}m_{\alpha}+\tilde{K}\varepsilon_{\alpha\beta}m_{\beta})+\eta_{\alpha}\;,$ (61)
@@
$C_{x}\left(\tau\right)=\frac{\langle x\left(t\right)x\left(t+\tau\right)\rangle_{t}}{\langle x\left(t\right)^{2}\rangle_{t}}.$ (62)
Note that $\langle x\rangle_{t}=0$. If this was not the case, one should replace $x(t)$ with $x(t)-\langle x\rangle_{t}$ in Eq. (62). As seen in Figure 16, even though the magnetization performs quite regular oscillations, its auto-correlation function decays, thereby indicating that this oscillator eventually forgets its initial phase [198, 129, 199, 200, 201, 202, 130, 203, 204, 205, 206, 207].
In many-body systems, however, interactions can overcome noise, producing coherent oscillations over arbitrarily long periods [127, 99, 102, 113, 208, 115, 117, 98, 97, 116, 209, 111, 112, 115, 114, 210, 208, 100, 211, 212, 95, 94]. Figure 16 shows that in the 3D nonreciprocal Ising model, for large system size, temporal correlations of individual spins $\sigma_{i}^{A}$ (panel e) and of the total magnetization $M_{A}$ (panel f) persist and do not decay over large times, making it a robust clock. This robust clock is an emergent feature of our model which does not rely on any periodic drive externally imposing a phase.
Fig. 16: Time correlations in the swap phase of the NR Ising model. (a) Time evolution of $m_{A}$ in the zero-dimensional mean-field equation with added noise. (b) Time evolution of a single spin $\sigma^{A}_{i}$ in the 3D NR Ising model. (c) Time evolution of the total magnetization of $A$-spins, $M_{A}$, in the 3D NR Ising model. (d-f) Auto-correlation function of $m_{A}$, $\sigma^{A}_{i}$, and $M_{A}$, respectively (defined in Eq. (62)). Panels a and d are obtained from numerical integration of Eq. (61) with the parameters: $\tilde{J}=1.2$, $\tilde{K}=0.12$, $\zeta=0.0009$, $dt=0.01$. Panels b-c and e-f are obtained from MC simulation with the parameters: $\tilde{J}=1.6$, $\tilde{K}=0.1$, $L=80$.
The different behaviors of the time correlations in the noisy clock Eq. (61) and in the 3D NR Ising can be respectively compared to spatial density correlations in a liquid (which decay with distance) and a solid (which is maintained over arbitrarily large distances). Such comparison makes Eq. (61) a “time liquid” and the 3D NR Ising a “time crystal” which has infinite coherence time.
Fig. 17: Diverging coherence time of the swap phase in the 3D NR Ising model. (a) The correlation function $C_{M_{A}}(\tau)$ of the total magnetization of $A$-spins, for different system sizes. Correlations persist over longer times as system size increases. (b) The envelope of the auto-correlation function, $\tilde{C}_{M_{A}}(\tau)$, shown for linear system sizes $L=12,16,20,24,28,32$ (from bottom to top). The coherence time $\tau_{c}$ (Eq. (63)) is calculated from the intersection of $\tilde{C}_{M_{A}}(\tau)$ with $e^{-1}$ (dashed black line). Inset: $\tau_{c}$ as a function of $L$ on a log-log scale. The dots are calculated from the data in the main panel and the line is the best linear fit, which has a slope of 3.01 with a standard error of 0.03. The coherence time is proportional to $L^{d}$, supporting a time-crystal behavior. Obtained from MC simulations with the coupling parameters $\tilde{J}=1.6$ and $\tilde{K}=0.1$.
To substantiate our claim, we find the coherence time of the swap phase in the 3D NR Ising model by determining the decay rate of the auto-correlation’s envelope, $\tilde{C}_{M_{A}}(\tau)$. We construct $\tilde{C}_{M_{A}}(\tau)$ by interpolating the peaks of $|C_{M_{A}}(\tau)|$. We then identify the coherence time $\tau_{c}$ as the time for which
@@
i.e. linearly with the number of spins, thereby proving the perfect time-crystal behavior of the 3D NR Ising in the thermodynamic limit. A similar scaling for $\tau_{c}$ has been observed in a periodically driven 2D system in Ref. [213].
This result suggests the existence of perfectly coherent oscillations in the thermodynamic limit $L\to\infty$, which can be seen as a breaking of ergodicity [214, 215, 216]. Even at arbitrarily long times, the probability distribution function $P(t)\equiv U(t)P_{0}$ does not necessarily converge to a constant stationary distribution $P^{\text{ss}}$ such that $U(t)P^{\text{ss}}=P^{\text{ss}}$. Here, $U(t)$ is the evolution operator acting on probability distributions, e.g. defined by the master equation (13). Instead, we expect that the system evolves towards a limit-cycle $P_{\text{LC}}$ satisfying $P_{\text{LC}}(t)=P_{\text{LC}}(t+T_{\rm osc})$ where $T_{\rm osc}$ is the oscillation period [217]. As a consequence, the distance between the distributions obtained from evolving two initial conditions $P_{1}$ and $P_{2}$ does not necessarily go to zero at long times, i.e. one can have
$\lim_{t\to\infty}\lVert U(t)P_{1}-U(t)P_{2}\rVert\neq 0$ (65)
because the many-body oscillator remembers its phase (i.e. its position on the limit cycle), which is encoded in the probability distribution.
We can combine Eq. (64) with the conjecture of Ref. [131] to obtain a bound on the minimum rate of entropy production required to maintain the oscillations. This conjecture bounds the average entropy per oscillation $\Delta S$ by the number of coherent oscillations $\mathcal{N}\equiv\tau_{c}/T_{\rm osc}$ through the inequality $\Delta S\geq 4\pi^{2}\mathcal{N}$. As a consequence, the rate of entropy production per unit volume
$\dot{s}\equiv\frac{\Delta S}{L^{d}T_{\rm osc}}$ (66)
must satisfy
@@
In the stable swap phase $T_{\rm osc}$ does not depend on $L$ (it converges to a constant as $L\to\infty$, see Fig. 24a in Sec. IX.1). As $\tau_{c}(L)\sim L^{d}$, this bound converges to a finite value when $L\to\infty$, meaning that there is a finite minimal cost per oscillator to maintain the time-crystal state. The fit in the inset of Fig. 17b provides $\tau_{c}/L^{d}$ for the example considered in the plot. In Appendix A we calculate the entropy production explicitly from simulations under the assumption of local detailed balance, and show that $\dot{s}$ converges to a finite value and that this value indeed exceeds the lower bound provided by Eq. (67).
Note that ordered oscillations and infinite coherence time are not decoupled properties. It is plausible that a locally coupled system can maintain spatially ordered oscillations over arbitrarily large distances only if all its parts adhere to the same prescribed phase determined by the initial conditions.
@@ VIII.3 The phase transition from disorder to swap in the high $(\tilde{J},\tilde{K})$ regime
Fig. 18: Destruction of the second-order phase transition due to scroll waves in the high $\tilde{K}$ and high $\tilde{J}$ regime. (a) $R$ vs. $\tilde{K}$ for $\tilde{J}=2$. Phase-space angular momentum is shown in the inset. (b) $\theta$ snapshot of a 2D slice from a 3D simulation with the parameters $\tilde{J}=2$, $\tilde{K}=0.9$, $L=320$, corresponding to the point marked by a black hollow circle on panel a. (c) The same snapshot coarse-grained over $10\times 10$ lattice sites. The snapshots show the destruction of order by scroll waves (appearing as spiral defects in the 2D cut). Obtained from MC simulations.
@@
At high $\tilde{J}$ and $\tilde{K}$ values, the transition from disorder to swap is different than for small $\tilde{J}$ and $\tilde{K}$ values. This is seen in Fig. 11 where the change in $R$ and $\mathcal{L}$ at the phase transition appears to become more abrupt in the upper right region, compared to the lower left region. We further analyze this transition in Fig. 18. Figure 18a shows the order parameters as a function of $\tilde{K}$, for fixed $\tilde{J}=2$, and various system sizes. The plots of $R$ and $\mathcal{L}$ vs. $\tilde{K}$ do not seem to converge to a second-order phase transition in which the order parameters are continuous. Instead, an abrupt change in $R$ and $\mathcal{L}$ emerges at high system sizes, suggesting a discontinuity as in first-order phase transitions. A clear critical point cannot be identified since convergence is not reached at $L=320$, the largest linear system size in our simulations (see $0.8\lesssim\tilde{K}\lesssim 1.1$ region in Fig. 18a).
@@
The sharp decrease in synchronization is caused by the emergence of highly noisy scroll waves that destructs the order. This is seen in Fig. 18b showing a 2D-slice snapshot of $\theta$ taken from the 3D simulation. Scroll waves are the 3D analog of spiral waves, having line defects that are typically closed into rings instead of point defects [142, 218, 86, 219] (see also Movie 3 and Fig. 27 in Sec. X in which the scroll waves are more visible). They also have spiral cross sections, which is seen in the 2D-cut $\theta$ snapshot in Fig. 18b and in its coarse-grained version, Fig. 18c (the latter figure makes the defects more apparent by averaging out the noise). We discuss scroll waves in more detail in Sec. X.
@@
The destruction of the swap phase by scroll waves is slightly reminiscent of the destruction of order in 2D due to spiral defects (compare Fig. 9e-f with Fig. 18a), except that here, it appears that the destruction only occurs for a limited range of high $\tilde{J}$ and $\tilde{K}$ values. Due to the size limitation of our simulations we cannot determine what portion of the 3D swap phase is destructed by the scroll waves or what kind of phase transition is associated with this destruction, although it appears to be first-order-like (in the sense that the order parameters are discontinuous).
VIII.4 Why is the swap phase stable in 3D?
@@
Fig. 19: Continuous symmetry breaking of the swap phase. (a) Comparison of discrete and continuous symmetry breaking in equilibrium and out-of-equilibrium models. (b) $M_{B}$ vs. $M_{A}$ for $\tilde{K}=0.5$, $\tilde{J}=2$, and $L=80$. i,ii, and iii show increasing numbers of MC sweeps. Each point is colored according to the sweep it corresponds to. Since the oscillation period is incommensurate, a continuous limit cycle is sampled in the infinite time limit. Hence, a continuous symmetry is broken. (c) Oscillation period of the magnetizations, $T_{\rm osc}$, in the 3D system as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and $L=80$. $T_{\rm osc}$ is measured as twice the average time between two subsequent sign flips of $M_{A}$. The continuous dependence of $T_{\rm osc}$ on the coupling parameter indicates incommensurate oscillations. (d) comparison between domain walls in 3D the Ising model (first row) and in the 3D NR Ising model (second row). At $t=0$, the two systems were in steady state and then perturbed by a domain wall. The domain wall is created by flipping the magnetization (Ising) and shifting the oscillation phase (NR Ising) for all spins located at $y>L/2$. System parameters are: $\tilde{J}=1.4$ (Ising) and $\tilde{J}=1.5$, $\tilde{K}=0.1$ (NR Ising), and $L=320$. i and iii are Snapshots of $\sigma$ (Ising) and $\theta$ (NR Ising, see color code in Fig. 2) in a 2D slice taken from a 3D simulation at $t=0$ and $t=1157$, respectively. ii Kymograph of $|{\rm d}{\bar{m}}/{\rm d}{y}|$ (Ising) and $|{\rm d}{\bar{\theta}}/{\rm d}{y}|$ (NR Ising), defined by the discrete symmetric derivative ${\rm d}f/{\rm d}y\equiv\left|\left(f_{y+5}-f_{y-5}\right)/10\right|$, where $\bar{m}$ is the magnetization and $\bar{\theta}$ is the phase, both averaged over the $xz$ plane. The smearing out of the dark pink region indicates the dissipation of the domain wall. Panels b-d are obtained from MC simulations.
@@
Our numerical evidence points to the existence of a stable swap phase. This relates to a longstanding question: can time-periodic states of noisy many-body systems with short-range interactions be stable in the thermodynamic limit [99, 102, 127, 113, 208, 115, 117, 98, 97, 116, 209, 111, 112, 115, 114, 210, 208, 100, 211, 212]. References [99, 102, 100] distinguished two cases, depending on whether the spontaneously broken time-translation invariance is discrete or continuous (Fig. 19a). In systems that evolve in discrete time as our Monte-Carlo simulation, this depends on the type of oscillations: commensurate vs. incommensurate. When the period of the oscillations is commensurate, i.e., a multiple of the discrete time-unite (a case referred to as “periodic $k$-cycle”), a discrete time-translation symmetry is broken. In that case, it was argued in [99] that the growth of out-of-phase droplets  [^4] Note that these potential droplets are not the droplets that destabilize the static-order phase (as in Fig. 21). Instead, these are droplets of a homogeneously oscillating state, immersed in an oscillatory background with a different phase destabilizes long-range ordered oscillations. That is because flat domain walls are expected to have a non-vanishing velocity, unlike in equilibrium [221], suggesting that a fluctuation producing a large enough droplet, can expand rather than shrink. A similar destabilization mechanism has been observed in discrete-symmetry flocks [222].
@@
In the incommensurate case, i.e. when the oscillation period is not a multiple of the discrete time-unite (e.g. the time step of the Monte-Carlo simulation), the discrete-time system is equivalent to continuous-time one (the discreteness is irrelevant). The oscillatory state then breaks a continuous time-translation symmetry, which can be seen by the fact that the oscillation’s phase at a given time can take any value between 00 and $2\pi$. In continuous symmetry breaking, domain walls smear out (as in the XY model) and thus the destabilization-by-droplets argument does not apply [223, 102, 100].
@@
In the nonreciprocal Ising model studied here, the oscillation period is incommensurate with the Monte-Carlo time step. This is seen both in Fig. 19b where a continuous limit cycle is formed in the infinite-time limit, and in Fig. 19c where the period of the oscillation, measured as twice the average time between two subsequent sign flips of $M_{A}$, is shown to change continuously with the interaction parameters (a continuously changing number is in general incommensurate with $1$). This means that, while stable oscillations are not guaranteed, they are not ruled out by the droplet argument in Ref. [99], see also Ref. [212]. Figure 19d further exemplifies this: it shows an analysis of 3D simulations of both the Ising model and the nonreciprocal Ising model with a flat domain wall introduced at location $y=0$ at time $t=0$ (the domain wall separates oppositely magnetizaed states for Ising, and two swap states at different phases for nonreciprocal Ising). As time evolves, the domain wall in the nonreciprocal Ising system is shown to widen and dissipate (on top of slightly moving in the $+\hat{y}$ direction), unlike the domain wall in the Ising system which remain steady. This suggests that there are no stable domain wall between oscillating states in the swap phase of the nonreciprocal Ising model.
The existence of a broken continuous symmetry (time-translation invariance) suggests a parallel with the Mermin-Wagner theorem [110, 82] in the time domain [109, 91]: continuous changes in the oscillation phase over space may behave as generalized Goldstone modes and their fluctuations preclude order in 2D yet allow for order in 3D and beyond. Note that this discussion disregards any potential phase transition to quasi-long range order in 2D that would resemble the Berezinskii-Kosterlitz-Thouless (BKT) transition in the XY model. While we have not proven the absence of such a phase transition in the nonreciprocal Ising model, we do not expect it to be present. The similarity with the XY model blurs when we depart from the Hopf bifurcation towards the high $\tilde{J}$ regime (where one might have hoped to observe a phase transition to quasi-long range order), since the discrete symmetry of the spins causes a SNIC bifurcation in mean-field and droplet nucleation in simulations (Sec. IX). Even the CGL equation with additive noise, which is rotationally symmetric and hence more related to the XY model than the NR Ising model, was suggested to not have a BKT transition  [106, 107, 177].
@@
An analogy between (de)synchronization and surface roughening provides further reasons to expect the swap phase to be stable in 3D and not in 2D. If we add random noise to the mean-field equation (30), it takes the form of a noisy reaction-diffusion system
@@
where $\vec{X}=x_{1},..,x_{n}$ and $D$ can be a matrix. Now assume that the noiseless and diffusion-less version of Eq. (68) has a limit cycle in steady state. To test whether the addition of diffusion and noise allows for synchronized oscillations, we assume that we are deep within the synchronized regime, where fluctuations from the global limit cycle are small and can be treated perturbatively. In this case, we can focus on the phase variable that describes the position of the system on the limit cycle, while neglecting other degrees of freedom. This is known as a phase reduction, see Refs. [85, 224, 225, 226, 227]. Under certain approximations, Eq. (68) can then be mapped to the Kardar–Parisi–Zhang (KPZ) equation [103]
$\frac{\partial\Psi}{\partial t}=\alpha\nabla^{2}\Psi+\beta\left(\nabla\Psi\right)^{2}+{\rm noise}$ (69)
where the deviation $\Psi$ of the local phase with respect to the phase of the deterministic limit cycle plays the role of the surface height in the KPZ equation (Fig. 20), see Refs. [85, 102, 106, 107, 228]. The KPZ equation predicts that interfaces are always rough in 2D, but that there is a phase transition between rough and smooth interfaces in 3D [229]. This suggests, by analogy, that a spatially coherent oscillating phase can exist in 3D, but not in 2D. Note, however, that the variable $\Psi$ in Eq. (69) is an angle (defined on a compact space, the circle $S^{1}$), so the results on the usual KPZ equation cannot be directly applied. In particular, the fact that the phase $\Psi$ is compact allows for the existence of topological defects, that may change the physics. We refer to Refs. [106, 107, 102, 230, 231, 232] for further details on the compact KPZ equation. In 3D, defect rings are expected to be small (at least for certain range of $\tilde{J}$ and $\tilde{K}$) and hence not to completely destroy long-range order [102].
@@
IX Destruction of the static-order phase via droplets